# Photonic Exceptional Points in Holography and QCD

### v4

@@
In this work, based on an analogy of holographic confining geometries and using complexified fields, we build the holographic toy model of third order photonic exceptional points (EPs) of ternary coupled microrings with gain and loss, which makes an open, non-Hermitian quantum system. In our model, we discuss the Ferrell-Glover-Tinkham sum rule for various combinations of gain and loss systems, and numerically find the behavior of spectra which matches with the experiments. We also discuss the inhomogeneous case of a holographic lattice for three-site photonic EPs. Additionally, in our holographic model, we numerically find the behavior of phase rigidity and the Petermann factor around EPs versus various parameters of the model. We also discuss the connections between recent developments in complexified, time-dependent entanglement entropy and EPs, and finally, we connect EPs and the $\theta$-vacuum of QCD through topological structures, partition functions, and winding numbers, and find a second-order EP in a perturbed $\theta$-vacuum model.
@@ 1 Introduction
Exceptional points (EPs) [25], as shown in figure 1, are branch-point singularities in the parameter space of a system where two or more eigenvalues and their corresponding eigenvectors merge and become degenerate. They are the boundaries between regions of unbroken PT-symmetry, where a Hamiltonian $H$ has all real eigenvalues, and regions of broken PT-symmetry where $H$ has at least one pair of complex conjugate eigenvalues.
@@
Here, we connect exceptional points and exceptional lines to the end-wall of holographic confining geometries, which can simulate the confinement/deconfinement phase transitions [27]. We suggest there is an interesting link here, as both EPs and the confining end-wall cause similar dramatic phase transitions, induce chaos, and cause chiral symmetry breaking. Also, it has been suggested that at EPs, the metric ceases to exist [28], similar to the effects of the end-wall in AdS/QCD models.
There are several other reasons which indicate such connection between AdS/QCD models and EPs. First, is the behavior of the quantum measure “fidelity.” In quantum information, one could use fidelity to find the exceptional points and their orders, as they only require ground-state calculations. The definition of fidelity is
@@
Another reason is that the non-Hermitian Dirac operators in QCD can be modeled by (higher-ordered) EPs. At finite quark chemical potential, the QCD Dirac operator, $D(\mu)=\gamma^{\mu}(\partial_{\mu}+igA_{\mu})+\mu\gamma^{0}$ becomes non-Hermitian, as its eigenvalues become complex and can exhibit spectral degeneracies. These degeneracies can coalesce and so the chiral phase transition in QCD can be approached via a spectrum collapse similar to an EP.
The similar behavior of chaotic systems near exceptional points and around the confining walls is another evidence. In [29], the chaos in the scatterings of open strings has been discussed, where they noticed that the bound-state creation in the process would leave its signatures on the plots of the amplitude $A_{4}$ versus the Mandelstam variable $s$. They noted that the peaks in the plots are more pronounced at small $s$, but they fade away at large $s$, which they interpreted as the effects of asymptotic freedom, where the gauge coupling becomes weaker, and therefore the binding energy of the bound states becomes smaller. This behavior can also be modeled by the non-Hermitian skin effect and its bound states inside a loop [30, 31]. Note that both in [29] and [30], a similar power-law decay has been found, which only matches with the non-Hermitian behavior, as it would be invisible for Hermitian baths.
Based on the above ideas and based on bottom-up holographic QCD, in this work, we build a toy model which can simulate third order EPs. First, we explain the optical setup of anti-PT symmetry consisting of ternary coupled optical resonators and microrings. Based on this system, we then build a holographic toy model to simulate EPs. We discuss various results coming from this model such as lasing spectra, eigenvalue trajectories, PT phase diagram, Ferrell-Glover-Tinkham (FGT) sum rule, transfer fraction to coherent peak and holographic Petermann factor and phase rigidities.
Then, we discuss the connections between exceptional points and timelike entanglement entropy and the behavior of Kirkwood-Dirac–type distribution in the presence of EPs.
Finally, we focus on QCD itself and first define a winding number for EPs based on the winding number in QCD and then we look for EPs in $\theta$-vacuum QCD. We actually could not find any EP in that model, however after adding a small angular coupling, we found a second-order EP there. We finish with a brief conclusion.
2 Photonic setup of anti-PT symmetry and higher-order exceptional point
@@
In the vicinity of a third-order exceptional point, and when the PT symmetry is being broken, as has been argued in [42], one can experimentally achieve a stable single-mode laser, which could then be used for different applications.
Figure 2: A ternary system, consisting of three coupled microrings. On the left, a PT-symmetric and on the right, an anti-PT-symmetric system with dissipative coupling are depicted. This setup with single-mode operation could lead to various orders of exceptional points.
@@
where $a_{n}$ are the energy amplitudes in the cavities, $\omega_{n}$ is the resonance frequency, $\lambda_{n}$ is the net gain/loss in microcavities, and $\kappa$ is the coupling factor between microresonators.
@@
The lasing spectra of symmetric and asymmetric gain–gain–gain and gain–loss–loss microrings, evenly pumped or partially pumped triple microrings, and unbroken or broken parity-time symmetric coupled microring lasers are shown in [42]. Now our goal is to build a holographic toy model to simulate their results and find a similar spectra, which we show in the next section.
3 Holographic confining model of coupled ternary resonators and EPs
In holographic QCD models, flavor branes are introduced to account for the presence of quarks. Similarly, each microresonator in the optical system can be viewed as a flavor brane, with its own dynamics influenced by the gain and loss perturbations. The coupling coefficients between the microresonators determine the interactions between the resonator modes. In the holographic dual, these coupling coefficients can correspond to bulk interactions between the flavor branes, affecting the meson spectrum and dynamics. Exceptional points in the optical system correspond to singularities in the eigenvalue spectrum, marking phase transitions between different symmetry phases. In the holographic dual, these exceptional points can be interpreted as critical points or the end-wall in the bulk geometry, associated with phase transitions in the gauge theory.
To construct the holographic dual, we can start with a 5-dimensional AdS spacetime with a metric that incorporates the effects of the flavor branes. The action governing the dynamics of the system includes terms for the bulk gauge fields, the flavor branes, and the interactions between them. By solving the equations of motion derived from this action, we can obtain the meson spectrum and study the phase transitions corresponding to the exceptional points in the optical system.
@@ 3 Holographic confining model of coupled ternary resonators and EPs
and $J_{\alpha}$ is the Bessel function. Here the spectrum of mass is $m^{2}=-k^{2}$. So in the free AdS geometry, the system is gapless and the spectrum is continuous, while in QCD the system is gapped and the spectrum is discrete, similar to the case near the EPs.
As for the boundary condition, one can assume that the wave function vanishes on the wall, i.e., $\psi_{k}(z_{0})=0$, or it becomes peaked at the wall, which lets us determine the spectrum by zeroes of the appropriate Bessel function [29]. The choice of the boundary condition then affects the structure of the amplitude only at low energies. For $M^{2}=0$ and $d=4$, the wave function is
@@
Instead of a hard wall, the AdS horizon then is cut softly with an exponential function $e^{-D}$. Then, the eigenvalues and eigenfunctions are [29]
@@
Here, $L_{n,\ell}(x)$ are the Laguerre polynomials. One should note that at EPs the norm of these wave functions vanishes and there is a square-root branch point in the interaction strength. When at the EP, the norm of a wave function vanishes, the two resonances coalesce, and the spectroscopic factors increase significantly. This effect has been used to model nuclei near the drip line [44], which specifically is an open system.
Another interesting top-down holographic model which can be used is Witten’s model with the metric
@@
where $f(U)=1-\frac{U_{\Lambda}^{3}}{U^{3}}$, and $U_{\Lambda}$ is the position of the wall, and $L^{3}=\pi g_{s}N_{c}\ell^{3}$. The coordinate $x_{4}$ has circular radius $R=\frac{3U_{\Lambda}^{1/2}}{2\pi L^{3/2}}$, setting the mass scale $M_{gb}=1/R$. The spectrum, i.e., the wave functions of the first three scalar modes of this model, are shown in figure 1 of [29], where all modes peak at $r=r_{0}$.
For this model, in [29], it has been shown that the spectra are not of a Regge form but mostly in the form of a KK-like spectrum as $m^{2}\approx aN^{2}+bN+c$, again similar to the $N$ levels coalescing of an N-dimensional matrix, where for complex symmetric matrices, $(N^{2}+N-2)/2$ parameters are needed to enforce the coalescence of $N$ levels [45]. So this point again suggests a connection between EPs and the holographic end-walls, as both show similar functional behaviors in $N$. This wall at $U_{\Lambda}$ also causes various phase transitions, changes in the entanglement entropy, mutual information, and other quantum information quantities, and induces chaos.
@@
where $\phi(z)$ is a dilaton-like background representing the effect of the microresonators’ gain/loss, which is analogous to the flavor branes’ backreaction. For a “soft-wall” model, a common choice is $\phi(z)=\kappa^{2}z^{2}$, where $\kappa$ sets a scale that can be tuned to reflect coupling strengths or exceptional points. The bulk action for a gauge field $A^{a}_{M}$, which represents mode dynamics of each resonator, and scalar $X$, representing symmetry-breaking effects, or the PT/anti-PT transition would be $S=S_{\text{bulk}}+S_{\text{flavor}}$.
@@
and $m_{5}^{2}$ is the $5D$ mass squared of the scalar field, which is related to the conformal dimension $\Delta$ of the dual operator as $m_{5}^{2}L^{2}=\Delta(\Delta-4)$.
@@
where $T_{3}$ is the brane tension, and $F_{MN}^{(i)}$ is the field strength on the $i$-th brane, each corresponding to a resonator. The coupling between branes, which represents the anti-PT symmetry and exceptional point interactions, can be included via off-diagonal terms in $F_{MN}$ or interaction terms in $X$.
@@
which controls the symmetry-breaking pattern, mapping to PT/anti-PT transitions.
So, the 3-mode microresonators are related to the 3 flavor branes or gauge fields $A_{M}^{(i)}$, the gain/loss is related to the background dilaton $\phi(z)$, and the coupling/anti-PT symmetry is related to the off-diagonal components in $F_{MN}$ or $X$. The exceptional points are related to the critical points in the bulk potential $X(z)$ or interaction term. The single-mode lasing would be related to the normalized lowest-energy mode in the bulk spectrum.
@@
where $\mathcal{M}(z)$ is the coupling matrix, representing gain/loss and anti-PT symmetry. For the anti-PT symmetry with ternary resonators, it would be
@@
One should note that these bulk equations are along the holographic direction $z$, and the off-diagonal couplings $\kappa_{ij}$ generate interactions between the branes. The imaginary entries $i\gamma$ encode the gain/loss, which realizes the anti-PT symmetry in the bulk. In the UV boundary, $z\to 0$, the Dirichlet boundary condition leads to $a_{i}(0)=a_{i}^{(0)}$, which corresponds to the driving/resonator amplitude input. In the IR boundary, $z=z_{IR}$, the normalizable solution $\partial_{z}a_{i}(z_{IR})=0$ selects the single-mode lasing.
@@
Therefore, starting from a $5d$ AdS metric with soft-wall, in the from of $ds^{2}=\frac{L^{2}}{z^{2}}e^{-\kappa^{2}z^{2}}(\eta_{\mu\nu}dx^{\mu}dx^{\nu}-dz^{2})$, we can find numerically three coupled bulk equations for flavor brane gauge fields $a_{1}(z),a_{2}(z),z_{3}(z)$, and the coupling matrix $\mathcal{M}(z)$ with gain/loss for anti-PT symmetry. We also can find the EP3 condition related to the triple degeneracy in $\mathcal{M}$ and also we can find the boundary conditions which select the physical modes.

### v5

@@
In this work, based on an analogy with holographic confining geometries and using complexified fields, we build a holographic toy model of third order photonic exceptional points (EPs) of ternary coupled microrings with gain and loss, which makes an open, non-Hermitian quantum system. In our model, we discuss the Ferrell-Glover-Tinkham sum rule for various combinations of gain and loss systems, and numerically find the behavior of spectra which matches with the experiments. We also discuss the inhomogeneous case of a holographic lattice for three-site photonic EPs. Additionally, we numerically find the behavior of phase rigidity and the Petermann factor around EPs versus various parameters of the model. We also discuss the connections between recent developments in complexified, time-dependent entanglement entropy and EPs, and finally, we connect EPs and the $\theta$-vacuum of QCD through topological structures, partition functions, and winding numbers, and find a second-order EP in a perturbed $\theta$-vacuum model.
@@ 1 Introduction
Exceptional points (EPs) [18], as shown in figure 1, are branch-point singularities in the parameter space of a system where two or more eigenvalues and their corresponding eigenvectors merge and become degenerate. They are the boundaries between regions of unbroken $\mathcal{PT}$-symmetry, where a Hamiltonian $H$ has all real eigenvalues, and regions of broken $\mathcal{PT}$-symmetry where $H$ has at least one pair of complex conjugate eigenvalues.
@@
Here, we make an analogy between exceptional points and exceptional lines with the end-wall of holographic confining geometries, which can simulate the confinement/deconfinement phase transitions [20]. We note that, as both EPs and the confining end-wall cause similar phase transitions, induce chaos, and cause chiral symmetry breaking, each one could modeled using the other one. Also, it has been suggested that at EPs, the metric ceases to exist [21], similar to the effects of the end-wall in AdS/QCD models.
There are several other reasons which indicate such connections between AdS/QCD models and EPs. First, is the behavior of the quantum measure “fidelity.” In quantum information, one could use fidelity to find the exceptional points and their orders, as they only require ground-state calculations. The definition of fidelity is
@@
Another reason is that the non-Hermitian Dirac operators in QCD can be modeled by (higher-ordered) EPs. At finite quark chemical potential, the QCD Dirac operator, $D(\mu)=\gamma^{\mu}(\partial_{\mu}+igA_{\mu})+\mu\gamma^{0}$ becomes non-Hermitian, as its eigenvalues become complex and can exhibit spectral degeneracies. These degeneracies can coalesce and so the chiral phase transition in QCD could be approached via a spectrum collapse similar to an EP.
The similar behavior of chaotic systems near exceptional points and around the confining walls is another evidence for this analogy. In [22], the chaos in the scatterings of open strings has been discussed, where they noticed that the bound-state creation in the process would leave its signatures on the plots of the amplitude $A_{4}$ versus the Mandelstam variable $s$. They noted that the peaks in the plots are more pronounced at small $s$, but they fade away at large $s$, which they interpreted as the effects of asymptotic freedom, where the gauge coupling becomes weaker, and therefore the binding energy of the bound states becomes smaller. This behavior can also be modeled by the non-Hermitian skin effect and its bound states inside a loop [23, 24]. Note that both in [22] and [23], a similar power-law decay has been found, which only matches with the non-Hermitian behavior, as it would be invisible for Hermitian baths.
Based on the above ideas and based on bottom-up holographic QCD, in this work, we build a toy model which can simulate third order EPs. First, we explain the optical setup of anti-$\mathcal{PT}$ symmetry consisting of ternary coupled optical resonators and microrings. Based on this system, we then build a holographic toy model to simulate EPs. We discuss various results coming from this model such as lasing spectra, eigenvalue trajectories, $\mathcal{PT}$ phase diagram, Ferrell-Glover-Tinkham (FGT) sum rule, transfer fraction to coherent peak and holographic Petermann factor and phase rigidities.
Then, we discuss the connections between exceptional points and timelike entanglement entropy and the behavior of Kirkwood-Dirac-type distribution in the presence of EPs.
Finally, we focus on QCD itself and first define a winding number for EPs based on the winding number in QCD and then we look for EPs in $\theta$-vacuum QCD. We actually could not find any EP in the pure case, however after adding a small angular coupling, we found a second-order EP there. We finish with a brief conclusion.
So we should emphasize that in sections 3 and 4, we build a bottom-up phenomenological, holographic-like toy model and present an approximate dictionary, while in section 5 we construct an analogy based on comparing two matrices for two different open quantum systems, where there is no holography argument in sections 5 and 6. Then, in section 7.1, we look for EPs in a perturbed or open $\theta$-vacuum QCD model directly (without any use of holography), first by a numerical search 7.1 and then analytically in 7.2.
2 Photonic setup of anti-$\mathcal{PT}$ symmetry and higher-order exceptional point
@@
In the vicinity of a third-order exceptional point, and when the $\mathcal{PT}$ symmetry is being broken, as has been argued in [25], one can experimentally achieve a stable single-mode laser, which could then be used for different applications.
Figure 2: A ternary system, consisting of three coupled microrings is shown. On the left, a $\mathcal{PT}$-symmetric and on the right, an anti-$\mathcal{PT}$-symmetric system with dissipative coupling are depicted. This setup with single-mode operation could lead to various orders of exceptional points.
@@
where $a_{n}$ are the energy amplitudes in the cavities, $\omega_{n}$ is the resonance frequency, $\lambda_{n}$ is the net gain/loss in microcavities, and $\kappa$ is the coupling factor between micro-resonators.
@@
The lasing spectra of symmetric and asymmetric gain-gain-gain and gain-loss-loss microrings, evenly pumped or partially pumped triple microrings, and unbroken or broken parity-time symmetric coupled microring lasers are shown in [25]. Now our goal is to build a holographic toy model to simulate their results and find a similar spectra, which we show in the next section.
3 Holographic confining toy model of coupled ternary resonators and EPs
In holographic QCD models, flavor branes are introduced to account for the presence of quarks. Similarly, each microresonator in the optical system can be viewed as a flavor brane, with its own dynamics influenced by the gain and loss perturbations. The coupling coefficients between the microresonators determine the interactions between the resonator modes. In a bottom-up, phenomenological holographic toy model, these coupling coefficients are related to bulk interactions between the flavor branes, affecting the meson spectrum and dynamics. Exceptional points in the optical system correspond to singularities in the eigenvalue spectrum, marking phase transitions between different symmetry phases. In a holographic dual perspective, these exceptional points can be interpreted as critical points or the end-wall in the bulk geometry, associated with phase transitions in the boundary side.
To construct a bottom-up, phenomenological holographic toy model, we can start with a 5-dimensional AdS spacetime with a metric that incorporates the effects of the flavor branes. The action governing the dynamics of the system includes terms for the bulk gauge fields, the flavor branes, and the interactions between them. By solving the equations of motion derived from this action, we can obtain the qualitative meson spectral and response features observed in non-Hermitian photonic systems with exceptional points, and then we can study the phase transitions corresponding to the presence of exceptional points in the optical system.
@@ 3 Holographic confining toy model of coupled ternary resonators and EPs
and $J_{\alpha}$ is the Bessel function. Here the spectrum of mass is $m^{2}=-k^{2}$. Note that in the free AdS geometry, the system is gapless and the spectrum is continuous, while in QCD the system is gapped and the spectrum is discrete, which is similar to the case near the EPs.
As for the boundary condition, one can assume that either the wave function vanishes on the wall, i.e., $\psi_{k}(z_{0})=0$, or it becomes peaked at the wall, which lets us determine the spectrum by zeroes of the appropriate Bessel function [22]. The choice of the boundary condition affects the structure of the amplitude only at low energies. For $M^{2}=0$ and $d=4$, the wave function is
@@
Instead of using a hard wall model, the AdS horizon then could be cut softly with an exponential function $e^{-D}$. Then, the eigenvalues and eigenfunctions would be [22]
@@
Here, $L_{n,\ell}(x)$ are the Laguerre polynomials. One should note that at EPs the norm of these wave functions vanishes and there is a square-root branch point in the interaction strength. When at the EP, the norm of a wave function vanishes, the two resonances coalesce, and the spectroscopic factors increase significantly. This effect has been used to model nuclei near the drip line [27], which specifically is an open system. Here we use such soft-wall models to construct our analogy.
Another interesting top-down holographic model which could be used in future works, is Witten’s model with the metric
@@
where $f(U)=1-\frac{U_{\Lambda}^{3}}{U^{3}}$, and $U_{\Lambda}$ is the position of the wall, and $L^{3}=\pi g_{s}N_{c}\ell^{3}$. The coordinate $x_{4}$ has circular radius $R=\frac{3U_{\Lambda}^{1/2}}{2\pi L^{3/2}}$, which sets the mass scale $M_{gb}=1/R$. The spectrum, i.e., the wave functions of the first three scalar modes of this model, are shown in figure 1 of [22], where all modes peak at $r=r_{0}$.
For this model, in [22], it has been shown that the spectra are not of a Regge form but mostly in the form of a KK-like spectrum as $m^{2}\approx aN^{2}+bN+c$, again similar to the $N$ levels coalescing of an N-dimensional matrix, where for complex symmetric matrices, $(N^{2}+N-2)/2$ parameters are needed to enforce the coalescence of $N$ levels [28]. So this point again suggests a connection between EPs and the holographic end-walls, as both show similar functional behaviors in $N$. This wall at $U_{\Lambda}$ also causes various phase transitions, drastic changes in the entanglement entropy, mutual information, and other quantum information quantities, and induces chaos.
@@
where $\phi(z)$ is a dilaton-like background simulating the effect of the microresonators’ gain/loss, which is analogous to the flavor branes’ backreaction. For a “soft-wall” model, a common choice is $\phi(z)=\kappa^{2}z^{2}$, where $\kappa$ sets a scale that can be tuned to reflect coupling strengths or exceptional points. The bulk action for a gauge field $A^{a}_{M}$, which simulate mode dynamics of each resonator, and scalar $X$, simulating the symmetry-breaking effects, or the $\mathcal{PT}$/anti-$\mathcal{PT}$ transition, would be $S=S_{\text{bulk}}+S_{\text{flavor}}$.
@@
and $m_{5}^{2}$ is the $5D$ mass squared of the scalar field, which is related to the conformal dimension $\Delta$ of the dual operator as $m_{5}^{2}L^{2}=\Delta(\Delta-4)$. Note that here the complex bulk parameters are only computational devices.
Note that here, the complex bulk couplings and effective masses are used as calculational devices to encode dissipative or gain-loss effects of the boundary theory. We are not claiming the existence of a fundamental non-Hermitian bulk gravitational theory. Rather, physical non-Hermiticity should be understood as entering through boundary conditions, external sources, or effective open-system descriptions, in line with Schwinger-Keldysh or influence-functional approaches.
@@
where $T_{3}$ is the brane tension, and $F_{MN}^{(i)}$ is the field strength on the $i$-th brane, each corresponding to a resonator. The coupling between branes, which represents the anti-$\mathcal{PT}$ symmetry and exceptional point interactions, can be included via off-diagonal terms in $F_{MN}$ or interaction terms in $X$.
@@
which controls the symmetry-breaking pattern, mapping to $\mathcal{PT}$/anti-$\mathcal{PT}$ transitions.
So, the 3-mode microresonators are related to the 3 flavor branes or gauge fields $A_{M}^{(i)}$, the gain/loss is related to the background dilaton $\phi(z)$, and the coupling/anti-$\mathcal{PT}$ symmetry is related to the off-diagonal components in $F_{MN}$ or $X$. The exceptional points are related to the critical points in the bulk potential $X(z)$ or interaction term. The single-mode lasing would be related to the normalized lowest-energy mode in the bulk spectrum.
@@
where $\mathcal{M}(z)$ is the coupling matrix, representing gain/loss and anti-$\mathcal{PT}$ symmetry. For the anti-$\mathcal{PT}$ symmetry with ternary resonators, it would be
@@
One should note that these bulk equations are along the holographic direction $z$, and the off-diagonal couplings $\kappa_{ij}$ generate interactions between the branes. The imaginary entries $i\gamma$ encode the gain/loss, which realizes the anti-$\mathcal{PT}$ symmetry in the bulk. In the UV boundary, $z\to 0$, the Dirichlet boundary condition leads to $a_{i}(0)=a_{i}^{(0)}$, which corresponds to the driving/resonator amplitude input. In the IR boundary, $z=z_{IR}$, the normalizable solution $\partial_{z}a_{i}(z_{IR})=0$ selects the single-mode lasing.
@@
Therefore, starting from a $5d$ AdS metric with soft-wall, in the from of $ds^{2}=\frac{L^{2}}{z^{2}}e^{-\kappa^{2}z^{2}}(\eta_{\mu\nu}dx^{\mu}dx^{\nu}-dz^{2})$, we can find numerically three coupled bulk equations for flavor brane gauge fields $a_{1}(z),a_{2}(z),z_{3}(z)$, and the coupling matrix $\mathcal{M}(z)$ with gain/loss for anti-$\mathcal{PT}$ symmetry. We also can find the EP3 condition related to the triple degeneracy in $\mathcal{M}$ and also we can find the boundary conditions which select the physical modes. It is important though to mention that the complex bulk fields here are just computational devices we use to simulate the spectral properties and they are not actually physical solutions.
@@
Note that in the UV boundary ($z\to 0$), the fields map to sources or VEVs of the dual operators, such as resonator amplitudes, and in the IR boundary ($z=z_{IR}$) one can impose normalizability or Dirichlet/Neumann condition, which corresponds to mode selection, such as single-mode operation. The dictionary for our specific setup of three-coupled microresonators is shown in table 1.
Table 1: Mapping to Optical System
   Optical System    Holographic Dual
   Microresonatores (3 modes)    3 Flavor branes/ gauge fields $A_{M}^{(i)}$
   Gain/Loss    Background dilaton $\varphi(z)$
   Coupling/anti-PT symmetry    Off-diagonal components in $F_{MN}$ or $X$
   Exceptional points    Critical points in bulk potential $X(z)$ or interaction term
   Single-mode lasing    Normalizable lowest-energy mode in bulk spectrum